### v1

# MuLan: Adapting Multilingual Diffusion Models for Hundreds of Languages with Negligible Cost

Abstract
In this work, we explore a cost-effective framework for multilingual image generation. We find that, unlike models tuned on high-quality images with multilingual annotations, leveraging text encoders pre-trained on widely available, noisy Internet image-text pairs significantly enhances data efficiency in text-to-image (T2I) generation across multiple languages. Based on this insight, we introduce MuLan, Multi-Language adapter,a lightweight language adapter with fewer than 20M parameters, trained alongside a frozen text encoder and image diffusion model. Compared to previous multilingual T2I models, this framework offers: (1) Cost efficiency. Using readily accessible English data and off-the-shelf multilingual text encoders minimizes the training cost; (2) High performance. Achieving comparable generation capabilities in over 110 languages with CLIP similarity scores nearly matching those in English (38.61 for English vs. 37.61 for other languages); and (3) Broad applicability. Seamlessly integrating with compatible community tools like LoRA, LCM, ControlNet, and IP-Adapter, expanding its potential use cases.
* equal contribution; ✉ corresponding author: wangwenhai@pjlab.org.cn.
Figure 1: Images generated by MuLan with different backbones, such as Dreamshaper 8, Dreamshaper XL Lightning, and Pixart-Alpha, using a variety of languages or mixed-language prompts.
1 Introduction
Recent diffusion models [11, 18, 19, 40, 44, 48] for content generation have attained stunning advancements in terms of both aesthetic quality and text-content alignment. However, these models still face substantial limitations in multilingual support. For instance, one of the most popular image-generation models, Stable Diffusion [33], and its successors [28, 11] only supports English and a few Latin-based languages. This language barrier restricts the model’s performance in multilingual contexts and hinders its applicability worldwide across diverse cultural and linguistic backgrounds.
The community primarily adopts two approaches to achieve multilingual text-to-image (T2I) generation. (1) The first is translation-based methods where input content is temporarily translated into English before generating the image. However, this approach often leads to inference delays, translation errors, and notable issues when handling slang or culturally nuanced content. (2) The other approach is native multilingual T2I models [48, 29, 1, 39, 19, 40], which is trained directly on high-quality images captioned in the target language. While this native approach improves image generation quality for non-English languages, it relies on extensive, carefully curated image-generation data in the target language, making it data- and resource-intensive. As a result, exploring a more efficient and generalizable approach to achieve strong multilingual generation capabilities remains challenging.
On the other hand, thanks to the development of computational power and dataset scale, many existing language models have achieved strong multilingual capabilities through training on large-scale internet data. For example, models trained on text data (e.g., BERT [8], the GPT series [2, 25], and LLaMA [42, 9]) and those trained on image-text pairs (e.g., CLIP [29], ALIGN [15], and InternVL [6]) demonstrate outstanding performance in multilingual understanding. Since the multilingual capability of T2I generation models is closely tied to their text encoders, it becomes essential to explore how these powerful multilingual text encoders can be leveraged to enable existing generative models to achieve multilingual capabilities more efficiently and effectively.
In this work, we deeply explore the application of multilingual semantic alignment in image generation from the perspective of language and image-text alignment. We also reveal that text encoders trained on large-scale multilingual image-text datasets with noisy data demonstrate remarkable data efficiency in multilingual image generation. Based on this insight, we introduce MuLan, a lightweight Multi-Language adapter that connects text encoders to diffusion models with low-cost training, facilitating native support for hundreds of languages. Specifically, we employ a plug-and-play language adapter with fewer than 20 million parameters to bridge a frozen multilingual text encoder with a frozen diffusion model. This setup exhibits outstanding zero-shot performance on multilingual T2I generation tasks, even relying on minimal English training data. Additionally, the framework is highly compatible with widely used community models and tools, such as LoRA [13], LCM [23], ControlNet [49], and IP-Adapter [47], demonstrating excellent flexibility, as shown in Figure 1.
Our model offers advantages in terms of training cost, multilingual performance, and adaptability compared to previous works. Specifically, MuLan leverages large-scale pre-trained multilingual text encoders, reducing the need for extensive multilingual datasets. As a result, the model can efficiently adapt to over 110 languages using only a small amount of English training data, achieving generation quality comparable to English (e.g., CLIP similarity scores: 38.61 for English, average 37.61 for other languages). Additionally, as shown in Figure 4, MuLan ’s plug-and-play architecture seamlessly integrates with existing model architectures and frameworks, greatly enhancing its compatibility with various community-driven tools and models.
In summary, our contribution is three folds:
(1) We investigate different possible ways to equip monolingual text-to-image models with multilingual ability, among which we demonstrate that using a text encoder properly trained from noisy web-scale data (such as InternVL [6]) is of great data efficiency.
(2) We propose MuLan, a plug-and-play lightweight language adapter that can be combined with any community models and tools for native multilingual generation in 110+ languages. Our method requires only English image-text pairs and avoids problems in previous attempts, including preparing datasets in a restricted number of languages, heavy computing budget, and inflexibility when combined with various community models/tools.
(3) We demonstrate our method’s effectiveness and efficiency through thorough quantitative and qualitative experiments. Notably, it requires only about 12 hours of training on 8 A100 GPUs and 17M English data to perform very well on 110+ languages, with very close CLIP similarity scores of 38.61 (English) v.s. 37.61 (average of other languages).
2 Related Work
Multilingual Diffusion Models.
Recent advancements have seen the rise of diffusion-based models, which have significantly improved image generation quality and diversity. Popular models such as Stable Diffusion series [33, 11, 28], DALL-E [32, 31], Imagen [35] and Glide [24] demonstrate photorealistic generation capabilities, yet they are primarily trained on English data and thus struggle with multilingual image generation. Although diffusion models using CLIP text encoders can generalize somewhat to Romance languages (e.g., French), their performance significantly degrades for languages outside this family, particularly East Asian languages such as Chinese. This limitation arises from the models’ reliance on English-centric text encoders, such as CLIP [29] and T5 [30], and the predominantly English training data.
Recent efforts have explored multilingual image generation by incorporating multilingual text encoders and datasets to overcome the limitations of English-centric models. One approach involves building models entirely from scratch with non-English data. For instance, Hunyuan-DiT [19] and Kolors [40] incorporate Chinese text encoders and extensive Chinese datasets to enhance their support for culturally specific concepts and improve generation quality in Chinese. Alternatively, some methods attempt to adapt existing models by replacing or fine-tuning the text encoder. Models such as Taiyi [44], PanGu [22], and AltDiffusion [46] replace the text encoder in Stable Diffusion and then fine-tune it with multilingual data, thus reducing the overhead compared to training from scratch. Despite these efforts, these models still face computational challenges, as fine-tuning often requires updating the entire denoising diffusion UNet.
Multilingual Language Model.
Language models have evolved through various modeling approaches, reflecting a range of training objectives and capabilities. Early models, such as BERT [8], BART [17], focus on understanding sentence structure and contextual relationships by predicting masked tokens within a sentence. More recent large language models (LLMs), such as LLaMA series [42, 9], T5 [30] and GPT series [2, 25], build on this foundation with improved context understanding and generation abilities, allowing them to handle a wide range of tasks with nuanced language comprehension and open-ended text generation. Building on these foundational text encoders, recent works, such as CLIP [29], ALIGN [15], and InternVL [6], incorporate visual alignment through paired image-text data, creating joint representations that bridge language and vision for cross-modal tasks.
Previous works such as mBERT [8] and LLaMA3 [9] have shown strong multilingual capabilities by pre-training on large multilingual corpora, enabling these models to understand and generate text in a variety of languages. InternVL-LLaMA [6] achieves powerful multilingual cross-modal capabilities by aligning a text encoder with a vision transformer (ViT) on the multilingual image-text dataset LAION[36], enabling effective image-text contrastive learning. However, due to the scarcity of large-scale multilingual image-text datasets, some approaches [5, 3] have resorted to using translated datasets and distillation learning to align multilingual text features. Despite these advances, an open challenge remains inefficiently leveraging multilingual text encoders for text-to-image (T2I) generation. Previous works, such as Taiyi [44] and AltDiffusion [46], have adopted multilingual text encoders and fine-tuned them on multilingual image-text pairs for T2I tasks. In contrast, our method takes advantage of InternVL-LLaMA’s multilingual capabilities, requiring only a small amount of English data and without the need to fine-tune the SD model weights, achieving state-of-the-art performance in multilingual T2I generation.
3 Proposed Method
In this section, we first revisit the mainstream text-to-image generation framework, analyzing its underlying structures and presently available data resources. Based on these insights, we introduce MuLan, a cost-effective multilingual generation framework designed to improve cross-language adaptability and generation quality.
3.1 Revisiting Text-to-Image Generation
Given an input text prompt $x$ and ground-truth image $y$ from the training dataset $D$, a mainstream text-to-image (T2I) generation model $G(\cdot)$, which consists of a language model $L(\cdot)$ and a visual generator $V(\cdot)$, can be defined as follows:
$\theta_{l}^{*},\theta_{v}^{*}=\arg\min_{\theta_{l},\theta_{v}}\mathbb{E}_{(x,y)\sim D}\left[\mathcal{L}(G(x;\ \theta_{l},\theta_{v}),y)\right],$ (1)
where $\theta_{l}$ and $\theta_{v}$ represent the parameters of the language model $L(\cdot)$ and the visual generator $V(\cdot)$, respectively, and $\mathcal{L}$ denotes the generation loss, * represents the optimal solution for function optimization. It can be observed that the primary components related to multilingual processing are the training dataset $D$ and the language model $L(\cdot)$. Therefore, in the following sections, we focus on the two modules, exploring pathways for building an efficient text-to-image generation model.
Text Encoder Architecture Supervision Method Aligned
BERT [8] encoder only MLM
T5 [30] encoder-decoder MLM
LLaMA [42, 9] decoder only NTP
CLIP-BERT [8] encoder only CL
InternVL-LLaMA [6] decoder-only NTP & CL
Table 1: Common Text Encoders. The common language model architectures and their supervision methods, “Aligned” indicates whether they have been aligned with images. In “Supervision Method”, “MLM” is Masked Language Modeling, “NTP” is Next Token Prediction, and “CL” is Contrastive Learning.
Language Model. Existing T2I generation models typically utilize pre-trained language models as text encoders. For instance, the Stable Diffusion series [33, 28, 11] utilizes CLIP text encoder [29] or T5 [30] as language encoders. Beyond these encoder-only and encoder-decoder language models, recent years have seen increased attention on decoder-only large language models (LLMs), such as the GPT series [2, 25] and LLaMA [42, 9]. These models are trained on pure text data by predicting the next word, supporting multiple languages, and outperforming traditional language models in NLP tasks. As shown in Table 1, a wide variety of language models are now available in the community; however, determining which models are suited for multilingual text-to-image generation remains an open question.
name num type lang-num quality
LAION-5B [36] 5B TI pairs 100+ Noisy
PixArt [4] 15M TI pairs 1 High
JourneyDB [26] 4M TI pairs 1 High
CCAligned [10] 100M Text Parallel 137 Noisy
CCMatrix [37] 69B Text Parallel 80+ Noisy
Table 2: Mainstream Dataset Types. The existing data used for T2I training mainly includes two formats: Text-Image pairs (TI pairs) and Text Parallel data.
Dataset. As shown in Table 2, the current datasets available in the community including multilingual text corpora, multilingual text-image paired datasets, and high-quality text-to-image datasets, hold potential value for multilingual image generation tasks. For example, large-scale datasets like LAION-400M/5B [36] contain large-scale noisy and lower-quality data, they provide multilingual text-image pairs, which are valuable for supporting multiple languages. Additionally, multilingual text translation corpora (such as CCMatrix [37]), while lacking corresponding image-text pair, offer useful cross-language correspondences. High-quality text-to-image datasets (such as JourneyDB [26]) provide high-resolution, high-quality images and are frequently used in text-image generation models; however, they primarily support English or some mainstream languages with limited coverage of other languages. Therefore, effectively utilizing low-cost, readily accessible internet-based text-image datasets and multilingual translation datasets to support multilingual text-to-image model training is a valuable area for further exploration.
3.2 MuLan: Toward Multilingual T2I Generation
Overall Architecture. To facilitate efficient and effective multilingual T2I generation, MuLan incorporates two key designs: (1) the multilingual semantic alignment through easily accessible large-scale data, and (2) a language adapter trained on a limited set of English T2I data. These designs enable MuLan to operate without the constraints of multilingual T2I data, allowing for more efficient training by leveraging existing models and datasets.
Multilingual Semantic Alignment. While many existing language models demonstrate robust multilingual capabilities, not all are well-suited for the multilingual T2I generation task needed for the Mulan framework. In this work, we emphasize the importance of maintaining a consistent vector space across languages for multilingual T2I generation. Specifically, given two text prompts, $x_{1}$ and $x_{2}$, that share the same meaning but are in different languages, and a text encoder $L(\cdot)$, the representations of these prompts, $L(x_{1})$ and $L(x_{2})$, should closely align. This alignment ensures that the conditional inputs to the image decoder remain consistent across languages, thereby preserving consistent image generation quality. Here, we mainly consider two alignment approaches: (1) Image-centered alignment;(2) Language-centered alignment.
Figure 2: Overview of of MuLan. The Transformer module is our Language Adapter, which follows the standard Transformer structure. We only train the Language Adapter, while all other modules are frozen. For different T2I model, we have designed different methods for using the Language Adapter.
(1) Language-Centered Alignment. A straightforward method that aligns multilingual semantics is to align multilingual semantics is to leverage a large set of translation data to align other languages’ vector spaces with the well-supported English vector space. By conducting distillation training with translation data alone, this can be achieved: we designate the Stable Diffusion language encoder as the teacher encoder, indicated by $L_{t}(x)$ and any multilingual encoder $L_{s}(x,\theta_{s})$ as the student encoder, aligning their features using MSE Loss. This alignment can be written as:
$\theta^{*}_{s}=\arg\min_{\theta_{s}}\mathbb{E}_{(x,y)\sim D_{\rm tr}}[{\rm MSE}(L_{s}(x_{1},\theta_{s}),L_{t}(x_{2}))]$ (2)
(2) Image-Centered alignment. In the image-centered alignment approach, CLIP maximizes the similarity between positive text-image pairs and minimizes it for negative pairs through contrastive learning. This training uses text-image pairs, and when the text includes multiple languages, the language encoder aligns different languages in the vector space naturally around the image. In this case, the objective function can be written as:
$\theta^{*}_{l}=\arg\min_{\theta_{l}}\mathbb{E}_{(x,y)\sim D}[{\rm cosine}(L(x,\theta_{l}),E_{I})],$ (3)
Here, $\theta_{l}$ refers to the parameters of the language model $L(\cdot)$, $E_{I}$ refers to the image feature. This method, however, is resource-intensive and requires large multilingual text-image pairs, which are challenging to obtain and require substantial storage. Using pre-trained models can help reduce these costs.
Language Adapter. After getting the aligned language model, to achieve cost-effective multilingual text-to-image generation, we propose MuLan. This model incorporates a lightweight language adapter $L^{\prime}$ that bridges a multilingual-aligned language model with a visual generator, enabling generalization to multiple languages after training on a small amount of English text-to-image generation data $D_{\rm en}$. So the Eqn 1 could be rewritten as:
$\theta^{*}_{l^{\prime}}=\arg\min_{\theta_{l^{\prime}}}\mathbb{E}_{(x,y)\sim D_{\rm en}}\left[\mathcal{L}(G(x;\theta_{l^{\prime}}),y)\right].$ (4)
After aligning different languages using Eqn 2 or Eqn 3, we can achieve multilingual T2I generation by training only this adapter. The key to low-cost implementation lies in freezing the language model and visual generator while training only the language adapter $\theta_{l^{\prime}}$ on small-scale English data $D_{\rm en}$.
In this work, we consider two types of adapters: MLP and transformer. These adapters are used to re-project the high-dimensional representations of text prompts from different languages into a unified space. We adopt different adapter designs for different diffusion models, as shown in Fig. 2. In detail, we can use either MLP or transfromer for projecting prompt embeddings for Pixart-$\alpha$ [4], and both architectures achieve good results. However, we find a simple MLP could not properly deal with Stable Diffusion models [33, 28]. Instead, we choose to use one layer encoder-decoder transformer with a set of learnable queries for extracting embeddings from InternVL outputs. For SDXL [28], we use two transformers to project embeddings for two text encoders that SDXL [28] adopts, and one attention pooling layer for extracting pool embeddings.
4 Experiments
This section describes the datasets, implementation details, and evaluation metrics used to assess our model, InternVL-MuLan. We compare its performance with state-of-the-art methods, including translation-based Stable Diffusion 1.5 [33] and AltDiffusion-m18 [46], across benchmarks such as XM18 [41], DPG-Bench [14], and COCO2014 [20]. Additionally, we perform ablation studies on adapter architectures and text encoders and present qualitative results to demonstrate our model’s capability to generate high-quality images across multiple languages.
4.1 Experimental setup
Datasets. We primarily use a subset of LAION-EN [36] with all samples that have aesthetic scores larger than 5.8 for training base models and PixArt [4] dataset for aesthetic models.
Implementation details. We trained MuLan adapters for a variety of pre-trained text encoders. By default, we use a transformer-based adapter with one encoder layer and one decoder layer. The number of transformer queries is set to 77 to match the input of Stable Diffusion [33].
All MuLan adapters are trained with AdamW [21] optimizer and 128 batch size. We use constant learning rate 1e-5 for Stable Diffusion 1.5/2.1 [33], 1e-6 for SDXL [28], and 2e-5 for Pixart-$\alpha$ [4]. For Stable Diffusion 1.5 [33], we train the adapter for 50k steps at the resolution of 512$\times$512, and for Stable Diffusion 2.1 [33] we adjust the resolution to 768$\times$768. For SDXL [28], we first train the adapter for 100k steps at the resolution of 512$\times$512 and finetune it for another 1k steps at the resolution of 1024$\times$1024. For Pixart-$\alpha$ [4], we train the adapter for 118k steps at the resolution of 512$\times$512. We randomly drop text conditions at the rate of 10% and use min-SNR [12] to accelerate training.
The training process was conducted on 8 NVIDIA A100-80G GPUs. For SD 1.5 and SD 2.1, training was completed within two days. For SDXL and Pixart-$\alpha$, training was completed within four days.
Evaluation Metrics. We use Crossmodal-3600 [41] to assess the model’s capabilities across 18 mainstream languages (denoted as XM18 hereafter). To evaluate the model’s generalization to additional languages, we tested multilingual versions of the COCO2014 [20] validation set and DPG-Bench [14]. We translated the prompts into 85 languages using Google Translate for these datasets. The model’s performance was compared with an ad-hoc translation-based SD1.5 model and other multilingual T2I models. Regarding evaluation metrics, we employed the Aesthetic Score to assess the quality of the generated images. Additionally, we calculated the Cosine Similarity (CLIP Sim) score with InternVL-LLaMA [6]. These metrics provide complementary insights into the visual quality and semantic accuracy of the model’s outputs.
4.2 Quantitive Results
Language InternVL-MuLan-SD15 Stable Diffusion (SD1.5)
AS CLIP Sim AS CLIP Sim
Arabic 6.33 38.22 6.11 34.43
English 6.35 38.61 6.06 39.57
French 6.29 37.96 6.05 38.16
German 6.24 37.95 6.01 38.37
Hindi 6.32 35.65 5.99 33.11
Italian 6.23 37.83 6.09 37.79
Japanese 6.27 37.96 6.02 36.67
Korean 6.33 37.57 6.06 36.36
Polish 6.32 37.07 6.09 36.69
Portuguese 6.32 37.04 6.08 37.19
Russian 6.33 37.89 6.02 37.38
Spanish 6.32 37.69 6.12 38.02
Thai 6.32 36.74 6.04 35.53
Turkish 6.35 37.97 6.09 37.68
Ukrainian 6.32 37.77 6.06 37.27
Vietnamese 6.25 37.94 6.04 37.20
Chinese 6.33 37.71 6.04 36.60
Dutch 6.36 37.41 6.12 37.59
Avg 6.31 37.61 6.06 36.97
Table 3: Comparison of zero-shot evaluation results between InternVL-Mulan-SD15 translation-based SD on XM18 [41]. InternVL-MuLan-SD15 consistently outperforms translation-based Stable Diffusion 1.5 in both aesthetic and CLIP similarity scores across most languages.
Language Model CLIP Sim
Chinese Taiyi-SDXL-3.5B [44] 35.44
Taiyi-SD-1B-Chinese [48] 36.84
InternVL-Mulan-SD15 37.84
InternVL-Mulan-SDXL 37.21
Japanese JapaneseSDXL [39] 38
InternVL-Mulan-SD15 38.01
InternVL-Mulan-SDXL 36.63
Table 4: Comparison of CLIP similarity scores on Chinese and Japanese with specialized models on XM18. InternVL-MuLan-SD15 surpasses dedicated Chinese and Japanese models despite not using Chinese or Japanese data during training.
Figure 3: Comparison of CLIP similarity scores of InternVL-MuLan-SD1.5 and AltDiffusion-m18 [46] across hundreds of languages on COCO2014 [20] val. Our model achieved performance comparable to AltDiffusion in mainstream languages while substantially surpassing AltDiffusion in less common languages.
Language Model Name Metrics
Overall Attribute Entity Global Other Relation
en SD1.5 62.66 74.02 72.95 72.11 74.95 75.28
AltDiffusion-m18 64.24 75.06 74.67 69.77 73.44 74.11
InternVL-MuLan-SD15 60.22 70.85 70.39 77.82 69.10 77.38
zh AltDiffusion-m18 62.18 72.33 73.58 71.22 72.55 74.30
InternVL-MuLan-SD15 58.75 71.04 71.33 66.44 71.74 70.23
ja AltDiffusion-m18 60.59 71.13 71.52 71.68 71.23 69.75
InternVL-MuLan-SD15 59.29 69.15 73.82 73.32 71.01 77.01
es AltDiffusion-m18 59.28 70.02 70.64 74.35 68.41 71.50
InternVL-MuLan-SD15 60.03 70.69 73.45 68.98 72.66 72.40
fr AltDiffusion-m18 59.41 69.81 69.51 72.59 69.27 70.38
InternVL-MuLan-SD15 60.83 72.74 71.99 73.11 71.06 75.34
Table 5: Performance metrics for different models across languages on DPG-Bench. InternVL-MuLan-SD15 achieves results close to those of AltDiffusion-m18 across various languages.
Multilingual T2I Comparison. We integrated InternVL-LLaMA [6] into the adapter’s model, which we call InternVL-MuLan. Specifically, for each version of Stable Diffusion, we train a separate adapter model, which we refer to as InternVL-MuLan-SD15, InternVL-MuLan-SD21, and InternVL-MuLan-SDXL.
We first compared the performance of InternVL-MuLan-SD15 with translation-based Stable Diffusion 1.5 [33]. The evaluation results on XM18 [41] are shown in Table 3. Our model significantly outperforms SD1.5 in terms of image aesthetic quality. For CLIP similarity scores, our model surpasses Stable Diffusion in most languages.
Next, we compared the performance of our model with AltDiffusion [46] in both mainstream and minority languages. We used DPG-Bench [14] to assess instruction-following ability in 5 mainstream languages. As shown in Table 5, our model’s instruction-following ability naturally extends to multiple languages, achieving results comparable to AltDiffusion [46]. To further evaluate our model’s capability in the broader range of minority languages, we evaluated our model on the COCO2014 [20] validation set (85 languages) and computed CLIP similarity scores, as shown in Figure 3(the complete list of scores will be provided in the appendix). Our model achieved performance comparable to AltDiffusion [46] in mainstream languages while substantially surpassing AltDiffusion [46] in less common languages.
Our model achieves state-of-the-art performance across multiple languages, reaching SOTA levels in mainstream languages and excelling in minority languages. Even though it was trained using only English text-image pairs, it outperforms certain specialized models in specific languages. While its performance in mainstream languages is slightly below that of AltDiffusion-m18 [46], the data and training costs we incurred are significantly lower than those of AltDiffusion. Specifically, our training on InternVL-MuLan-SD15 required only $0.5\times 8$ GPU-days, compared to the $19\times 64$ GPU-days needed by AltDiffusion-m18, highlighting a substantial difference in resource consumption. This makes our method both highly efficient and effective.
Figure 4: Examples of our model integrates with community tools. MuLan can be equipped with MVDream [38] to generate 3D models, and it is also directly compatible with existing community tools such as LoRA [13], ControlNet [49], and LCM [23].
Adapter Type #params CLIP
MLP 30M
1-layer Transformer 18M 38.6
4-layer Transformer 67M 38.3
Table 6: Comparison of Different Adapter Types for InternVL-MuLan-SD15. The 4-layer transformer refers to a Transformer adapter with 4 encoder layers and 4 decoder layers. CLIP similarity scores are reported on XM18 [41]. ✗means the model generates images of extremely low quality.
Ablation Study on Adapter Architecture. We studied the impact of adapter types and the number of transformer layers on InternVL-LLaMA [6]. The results are shown in Table 6. The results show that the Transformer adapter outperforms the MLP adapter under different text encoder choices. The possible reason is that synonymous prompts have different word orders in different languages, and the Transformer adapter can possess the representational capabilities of a language model.
Comparison of Different Text Encoders. We trained Transformer adapters with frozen SD 1.5 on BERT [8], T5 [30], LLaMA [42, 9], and their image-centered multilingual alignment versions (if available).
Model Align Method en es fr zh
LLaMA2-7B [42]
mT5-xl [45]
MultiLang-CLIP [3] Language-centered 35.8 33.6 34.2 31.8
AltClip-m18 [5] Language-centered 35.2 33.9 34.2 31.9
ViT-XLM-R-L* [7] Image-centered 36.5 36.7 37.2 34.5
InternVL-LLaMA [6] Image-centered 38.6 37.7 38.0 37.7
Table 7: Text Encoder Clip Similarity Scores Across Different Languages. ViT-XLM-R-L* stands for CLIP-ViT-H-14-frozen-xlm-roberta-large-laion5B-s13B-b90k [16].
The results are shown in Table 7. We found that for multilingual text encoders trained purely on language, such as XLM-RoBERTa-large [7] and LLaMA [42, 9], the performance is poor when training an adapter with only English text-image pairs. They are unable to generate coherent images when we freeze these text encoders. Moreover, they cannot generalize to other languages through training on English text-image pairs. However, when we train these text encoders using the image-centered approach or align them to CLIP text encoder [29] using the language-centered method, MuLan can connect them to Stable Diffusion [33], enabling the T2I model to achieve multilingual image generation capability.
4.3 Qualitative Results
Our model can generate high-quality images across multiple languages. It also supports multilingual mixed input and can recognize certain emojis. In addition to simple T2I tasks, there are numerous downstream applications of Stable Diffusion (SD) within the community, including LoRA [13], LCM [23], and ControlNet [49]. These tools played crucial roles in enhancing model adaptability, control over outputs, and finetuning for specific tasks. Unlike AltDiffusion [46], our model does not require finetuning on SD, allowing it to seamlessly integrate and be compatible with these community-developed SD applications in a plug-and-play manner. We show some examples in Figure 4.
5 Conclusion
We introduce language adapter MuLan that could equip image/video/3D diffusion models with multilingual generation abilities. MuLan shows strong zero-shot capabilities for up to 110 different languages, even if the adapter is solely trained on English data. MuLan also can be trained with a frozen text encoder and diffusion denoising model, which makes it applicable for many downstream models, such as LoRA [13], ControlNet [49], LCM [23], and etc., without any additional finetuning. MuLan is currently trained with paired data, and it could inevitably bring in bias and cause a distribution shift of original models. A promising extension would be to alleviate the need for paired data and make original capabilities intact. Furthermore, MuLan currently focuses on improving multilingual generation capabilities, but it would be interesting to extend it to improving prompt understanding and following under a multilingual context.
[…]
Supplementary Material
Appendix A More Experiments
In this section, we provide supplementary analyses to further evaluate and understand the performance and efficiency of our proposed model. These studies delve into key aspects such as training efficiency, adapter architecture, semantic alignment methods, and the model’s generalization across multiple languages. By examining the impact of various design choices and alignment strategies, we aim to shed light on the model’s multilingual capabilities and resource efficiency.
A.1 Training Efficiency
Our model leverages the multilingual capabilities of the InternVL-LLaMA text encoder, enabling effective training with minimal cost. In this section, we examine the changes in the model’s performance when data volume and training iterations are reduced.
$\lambda$ of samples Mean CLIP Sim
1 17M 35.80
$1/2$ 8.5M 35.46
$1/4$ 4.25M 35.40
$1/8$ 2.1M 35.23
$1/16$ 1M 35.17
$1/64$ 250k 35.13
$1/256$ 63k 35.08
$1/1024$ 16k 33.49
Table 8: Impact of dataset size on multilingual performance of the model. We evaluated the model’s performance using different proportions of the original dataset, ranging from full size ($\lambda=1$) to $1/1024$, and measured the Mean CLIP similarity scores across 7 languages. The results show that the model maintains relatively strong performance even with significantly reduced data sizes, demonstrating its efficiency and robustness under resource-constrained settings.
Setting. We choose Stable Diffusion 1.5 [33] as the backbone model, using the AdamW optimizer for training. The model was trained to convergence with a learning rate of 2e-5 and a batch size of 128, without employing any training tricks. Additionally, the dataset size was reduced, and we set multiple data proportion levels ranging from 0.5 to 0.001 to evaluate the impact of data size on the model’s performance.
Results. As shown in Table 8, we found that the model’s multilingual performance decreased as the data volume was reduced; however, it still maintained a relatively strong performance until we reduced the data to $1/1024$ of the default data volume. This level of data volume and training cost is highly developer-friendly, requiring only 48 GPU hours to achieve decent multilingual text-to-image (T2I) capabilities for the model. Compared to existing multilingual T2I models, such as AltDiffusion [46], our approach requires significantly less data and computational cost at every stage.
A.2 PixArt-$\alpha$ with MuLan
We also trained MuLan combined with Pixart-$\alpha$, which we called InternVL-MuLan-Pixart-$\alpha$. We evaluated InternVL-MuLan-Pixart-$\alpha$ on the Crossmodal3600 dataset, and the model demonstrated strong multilingual text-to-image (T2I) capabilities, performing exceptionally well in common languages. This also confirms that our approach is effective not only for Diffusion models with a UNet [34] architecture but also for models based on the DiT [27] architecture.
en ar ja zh fr
Ours 39.52 39.12 40.38 39.28 40.17
AD18 39.98 38.33 39.37 38.98 40.01
Table 9: CLIP similarity scores of InternVL-MuLan-Pixart-$\alpha$ and AD18 (AltDiffusion-m18 [46]) on crossmodal3600 [41]. Our model outperforms across multiple languages.
A.3 Comparison of Semantic Alignment Methods
Here, we conducted a comparison of two semantic alignment methods to explore the impact of different semantic alignment approaches on multilingual image generation.
Setting. We selected XLM-RoBERTa-Large[7] as the text encoder for comparison.
(1) For language-centered alignment, we conducted distillation training following Multilang-CLIP [3]. We utilized a larger translation dataset, CCMatrix [37], consisting of 2.91B samples. The choice of a large translation dataset was made to match the scale of LAION-5B [36], enabling a comparison between the effectiveness of translation data and image-text data in aligning multilingual representations. We selected CLIP text encoder as the teacher model and XLM-RoBERTa-Large [7] as the student model. The training process is configured with a batch size of 512 and a constant learning rate of 2e-5, applied over 10 epochs.
(2) For image-centered alignment, we selected ViT-XLM-R-L [16] provided by LAION and InternVL-LLaMA [6] as the text encoders. These text encoders were trained on the LAION-5B [36] dataset and aligned with ViT using contrastive learning.
Model Align Method en es fr zh
LLaMA2-7B [42]
XLM-R-L [7]
CCMatrix XLM-R-L [3] Language-centered 36.3 34.2 35.1 33.2
LAION5B XLM-R-L [16] Image-centered 36.5 36.7 37.2 34.5
InternVL-LLaMA [6] Image-centered 38.6 37.7 38.0 37.7
Table 10: Text Encoder Clip Similarity Scores Across Different Languages. CCMatrix XLM-R-L refers to XLM-RoBERTa-large pre-trained using CCMatrix [37] dataset with distillation learning. LAION5B XLM-R-L refers to CLIP-ViT-H-14-frozen-xlm-roberta-large-laion5B-s13B-b90k [16].
Results. The results are shown in Table 10. We can observe that although translation data is easier to collect and more efficient for training compared to text-image pairs, the performance achieved through text-centered alignment using translation data is inferior to that of image-centered alignment using text-image pairs. This highlights the critical importance of text-image pairs in image generation tasks.
Considering XLM-RoBERTa-Large [7] and other unaligned text encoders, it is evident that under our default experimental setup, training a single language adapter alone is ineffective. Even when trained on English text-image pairs, the model performs poorly when using English prompts and fails to generalize to other languages. However, after the model is aligned through either language-centered or image-centered approaches, it can gain multilingual text-to-image generation capabilities via MuLan.
Id SD15 SD21 SDXL AD18 Id SD15 SD21 SDXL AD18 Id SD15 SD21 SDXL AD18
af 35.39 34.54 34.68 33.83 id 35.67 35.55 35.49 31.53 ps 30.74 30.87 31.41 28.61
ar 39.00 38.20 38.13 38.77 ig 30.65 30.80 32.10 30.09 ro 37.17 36.88 36.93 30.51
az 34.19 33.67 33.66 30.95 is 30.68 30.96 31.97 29.17 ru 38.11 37.86 37.63 38.51
be 35.24 35.23 35.20 30.74 it 37.13 36.62 36.56 37.62 sd 30.61 30.20 30.26 27.93
bg 38.76 38.40 38.38 33.60 iw 38.52 38.07 38.17 26.84 sk 36.12 36.12 36.02 29.09
bn 33.58 32.56 33.61 30.06 ja 38.05 38.14 37.82 39.43 sl 36.83 36.28 36.22 27.21
bs 37.20 36.89 36.86 28.10 ka 35.70 35.53 35.80 30.38 sm 29.94 30.24 32.44 27.94
ca 37.74 37.43 37.22 33.17 kk 32.18 31.70 31.44 28.75 sn 29.61 29.71 30.74 29.97
co 35.03 34.50 34.39 33.95 km 25.82 26.71 28.43 24.81 so 29.67 29.05 31.38 29.05
cs 35.30 34.85 34.71 28.78 kn 28.94 27.42 29.65 28.59 sq 35.08 34.41 34.77 29.06
da 36.25 35.99 35.77 31.78 ko 38.06 37.21 37.09 38.27 sr 37.77 37.90 37.91 31.09
de 36.79 36.98 36.87 38.16 ku 29.86 29.23 30.41 27.50 st 29.76 30.00 32.24 29.63
el 35.39 35.49 35.24 25.36 ky 32.49 32.08 32.37 30.33 su 32.65 32.00 32.91 30.62
eo 31.17 30.81 30.99 28.92 la 30.25 30.32 30.39 29.71 sv 35.91 36.22 36.07 31.83
es 37.38 37.41 37.37 38.41 lb 30.96 30.59 30.93 30.28 sw 30.39 30.30 30.77 32.23
et 34.43 34.85 34.76 28.22 lt 36.02 35.94 35.77 27.00 ta 31.00 29.69 32.20 32.47
fa 38.22 37.99 37.92 28.58 lv 35.08 34.91 34.56 27.03 te 28.07 27.22 29.82 27.21
fi 37.16 37.16 36.74 28.71 mi 30.34 30.60 32.75 29.54 tg 30.84 30.43 31.29 28.34
fr 37.38 37.57 37.41 38.81 mk 38.17 38.29 38.41 31.66 th 36.95 36.83 36.76 37.56
fy 32.73 31.97 31.90 31.64 ml 31.75 29.67 32.78 29.72 tr 35.89 35.65 35.57 36.17
ga 27.56 26.96 28.26 28.27 mn 33.39 33.91 34.63 30.04 uk 37.62 37.22 36.98 37.94
gd 27.95 27.52 28.69 28.92 mr 35.14 34.12 34.15 31.52 ur 34.82 34.30 34.10 29.01
gl 37.06 36.80 36.59 36.08 ms 35.25 35.04 35.01 30.61 uz 30.18 29.50 30.24 28.91
gu 28.25 27.83 29.09 31.06 mt 29.97 29.22 30.64 29.87 vi 38.13 37.67 37.60 38.31
haw 31.31 31.52 33.81 31.37 my 28.15 28.10 29.56 27.90 yi 29.22 29.01 29.49 27.86
hi 36.94 36.47 36.28 36.92 ne 34.37 33.53 34.34 32.72 yo 30.32 29.95 31.68 28.65
hr 37.25 36.94 36.93 28.00 nl 36.85 36.64 36.39 37.91 zh 38.85 39.26 39.50 40.30
hu 36.38 36.08 35.98 26.44 no 35.94 35.80 35.73 31.65
hy 33.36 32.66 32.98 26.73 pl 35.97 35.82 35.30 37.12
Table 11: CLIP Sim on the COCO2014 validation set. ‘Id’ indicates the key of different language. SD15/21/XL [33, 28] represent our model implemented on the corresponding three T2I backbones, while AD18 refers to the AltDiffusion-m18 [46]. Our model also demonstrates strong text-to-image (T2I) capabilities across a wider range of languages.
A.4 Detailed Results on COCO2014 validation set
In Section 4.2, we evaluated InternVL-Mulan-SD15 on the COCO2014 [20] validation set (85 languages) and compared it with AltDiffusion-m18 [46]. More results are shown in Table 11. It can be observed that our model generalizes well to a wider range of languages and delivers impressive performance.
A.5 Impact of Semantic Alignment on Multilingual Features
In this section, we conduct a preliminary analysis of the multilingual output features produced by various text encoders to reveal the effects of the two image alignment training methods introduced in Section 3.2. The text encoders we selected for this analysis include XLM-RoBERTa-Large [7], CCMatrix pre-trained XLM-RoBERTa-Large in A.3, ViT-XLM-R-L* (LAION-5B [36] pre-trained XLM-RoBERTa-Large [7]), LLaMA2-7B [42], and InternVL-LLaMA [6].
We employ t-SNE, a nonlinear dimensionality reduction technique, to visualize the aggregation effects of multilingual features produced by these text encoders. t-SNE is particularly suited for preserving the local structure of high-dimensional data in a low-dimensional space. Ideally, for text encoders trained with multilingual alignment methods, the features corresponding to semantically equivalent prompts in different languages should exhibit aggregation effects when projected into a lower-dimensional space.
For this analysis, we randomly sampled 20 captions from the COCO2014 [20] validation set and translated them into 8 languages using machine translation, resulting in a total of 160 textual inputs. These inputs were then encoded using the selected text encoders to obtain their corresponding pooled feature representations.
Figure 5: t-SNE analysis on embeddings of 9 prompts in 20 languages produced by 5 text encoders. (a) XLM-RoBERTa-Large [7] (b) CCMatrix pre-trained XLM-RoBERTa-Large [7] in A.3 (c) Laion-5B [36] pre-trained XLM-RoBERTa-Large [7] (d) LLaMA2-7B [42] (e) InternVL-LLaMA [6]. Points of the same color represent embeddings corresponding to the same prompt translated into different languages.
The results are shown in Figure 5. As shown in Figures (a) and (d), the sample points in the t-SNE [43] plots exhibit a scattered distribution for text encoders that have not undergone image alignment training. This indicates that although XLM-RoBERTa [7] and LLaMA2 [42] possess multilingual representation capabilities, their output features for synonymous multilingual prompts are not closely located in the Euclidean space. In contrast, as shown in Figures (b)(c)(e), the sample points exhibit a clustered distribution after alignment training with images, with each cluster corresponding to the 20 different language versions of a single prompt. Through alignment training with images, the model minimizes the semantic discrepancies between multilingual representations, ensuring that the embeddings of synonymous prompts converge in the vector space.
Appendix B More Qualitative Results
In this section, we present more examples of multilingual generation by the model, as well as examples of its interaction with existing community models and tools.
B.1 Robustness to Multiple Languages
As shown in Figure 7, our model supports multiple languages, allows prompt inputs that combine different languages, and even recognizes emojis. For example, We can use the car emoji as a prompt (the first image in Figure 7), and the model can generate an image of a car.
B.2 Plug and Play on Different Visual Generator
Our model can be seamlessly integrated into existing fine-tuned models, such as DreamShaper, Realistic Vision, and others. In Figure 7, we present additional examples, all generated by existing models interacting with Mulan. These examples use prompts in multiple languages, demonstrating support for various language combinations as inputs.
Figure 6: Multilingual generation results. Our model supports multilingual and mixed-language inputs.
Figure 7: IP-Adapter Results. Our model enables multilingual style transfer by integrating with the IP-Adapter [47].
Our model also supports some existing tools based on the Stable Diffusion series developed by the community. Here, we showcase several popular models, including LoRA [13] models, ControlNet [49] models, and IP-Adapter [47] models. The model’s multilingual capabilities naturally come into play in these applications.
Figure 8: Lora Results. Our model can naturally support multilingual input when using LoRA [13].
Figure 9: ControlNet Results. Our model can utilize existing ControlNet [49] models, enabling multilingual image generation with conditional inputs such as depth maps and pose images.